Imagine that 11 people take a 10 question test. Suppose one student gets 10 correct answers, one gets 9  correct, one gets 8 correct, and so on, all the way down to the student getting none correct. What is the mean score, accurate to three decimal places?


1
Expert's answer
2021-04-14T07:27:34-0400

10 + 9 + 8 + 7 + 6 + 5 + 4 + 3 + 2 + 1 + 0 = 55

Mean score = 55 / 11 = 5

The mean score in three decimal way would be 5.000.